The product of 2 consecutive odd numbers is 63. What is the larger number?
Answer by stanbon(75887)

The product of 2 consecutive odd numbers is 63. What is the larger number?
1st: 2x-1
2nd: 2x+1
-------------------------
Equation:
(2x-1)(2x+1) = 63
4x^2-1 = 63
4x^2 = 64
x^2 = 16
x = 4 or -4
---
If x = 4
1st: 2*4-1 = 7
2nd: 2*4+1 = 9
--------------------
If x = -4
1st: 2x-1 = -9
2nd: 2x+1 = -7
======================
Cheers,
Stan H.